OpenStudy (anonymous):

evaluate g(x)=2x+5 for g(9)

hero (hero):

g(x) = 2(x) + 5 g(9) = 2(9) + 5

OpenStudy (anonymous):

You substitute 9 as the input into the function and solve

OpenStudy (skullpatrol):



hero (hero):

Guys, I believe she figured it out already.

OpenStudy (skullpatrol):

Thanks to you :-)